OpenStudy (anonymous):

How do I get from this step sqrt3(sqrt3/2sin(x)+1/2cos(x))-(sqrt3/2cos(x)-172sin(x)/4cos(x)(sqrt3/2sin(x)+172cos(x))-4sin(x)(sqrt3/2cos(x)-1/2sin(x)

OpenStudy (anonymous):

to this step 2sinx/2cos²x+2sin²x

OpenStudy (anonymous):

please use equation function

OpenStudy (anonymous):

\[\sqrt3(\sqrt3/2\sin(x)+1/2\cos(x))-(\sqrt3/2\cos(x)-172\sin(x)/4\cos(x)(\sqrt3/2\sin(x)+172\cos(x))-4\sin(x)(\sqrt3/2\cos(x)-1/2\sin(x)\]